(b) A representative consumer has a utility function U (x, y) = xy. A representative firm makes

good x and has a production function x = f(k, l) = (kl)0.25and an unavoidable fixed cost

equal to A. There are 100 consumers and, initially, 100 firms. Prices are w = v = Py = 1 and Px is determined in a competitive market. Representative consumer income is I = 2.

(b) Find the representative firm’s supply of good x?
Answer #1

production Here the fun of good x is given by - -3/4 114 MPL = 1/4L K x = K L - - 3/4 144 MPK = 1/4K L and - MRTS = MPL / M